# Distributional Lusternik–Schnirelmann category of manifolds

### v2

@@ 5.B. $4$-manifolds with free fundamental group
We will see more computations for $\operatorname{\mathsf{dcat}}$ of closed $4$-manifolds in Section 6.C.
5.C. An analog of Rudyak’s conjecture
@@
In the above setting, if $N={\mathbb{R}}P^{n}$ for odd $n$, then Question 5.12 is answered since $M$ is not contractible and thus, $\operatorname{\mathsf{dcat}}(M)\geq 1=\operatorname{\mathsf{dcat}}({\mathbb{R}}P^{n})$.
For $n\leq 2$, $\operatorname{\mathsf{dcat}}$ and $\operatorname{\mathsf{cat}}$ of closed orientable $n$-manifolds coincide, [DJ, Section 6.1]. So, Question 5.12 is answered affirmatively for $n\leq 2$ in view of Theorem 5.11.
@@
The proof of Theorem 5.11 for $n\in\{3,4\}$ uses the notions of category weight and detecting elements (see [R1, Section 2]). Unfortunately, we do not have useful analogous notions in the distributional theory yet, so we attempt to answer Question 5.12 affirmatively in some cases using our results from this paper.
@@
In case (1), we get $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)$ due to Theorem B (see Section 5.A). Similarly, in cases (2) and (3), $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)$ by our discussions in Section 5.B. In case (4), we apply Theorem 4.1 in dimensions $3$ and $4$ to get $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)$ again. Therefore, in all these cases, we use Theorem 5.11 to obtain the inequality $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)\geq\operatorname{\mathsf{cat}}(N)\geq\operatorname{\mathsf{dcat}}(N)$. ∎
@@ 6. For generalized connected sums
6.A. Sums along submanifolds
Let $M$ and $N$ be two closed smooth orientable $n$-manifolds for $n\geq 3$. Let $S$ be a closed orientable manifold of dimension at most $n$. Suppose there are embeddings $i:S\to M$ and $j:S\to N$ with respective normal bundles $\mu$ and $\nu$ and an orientation-reversing bundle isomorphism $g:\mu\to\nu$. Such a map $g$ exists, in particular, when $\dim(S)=n-2$ and $e(\mu)=-e(\nu)$ for the Euler characteristics. Using smoothness, we can choose tubular neighborhoods $U\subset M$ and $V\subset N$ of $i(S)$ and $j(S)$, respectively, such that $U$ and $V$ are identified with the total spaces of $\mu$ and $\nu$, respectively. Then $g$ yields a diffeomorphism $h:U\to V$ that maps $i(S)$ to $j(S)$. This produces an orientation-reversing diffeomorphism $\phi:U-i(S)\to V-j(S)$, see [IKRT, Section 3]. Let us denote this homeomorphic space by $L$. We take the disjoint union $(M-i(S))\sqcup(N-j(S))$ and identify $U-i(S)$ with $V-j(S)$ using $\phi$ to obtain a closed orientable $n$-manifold $M\#_{L}\hskip 0.56905ptN$, which we call the generalized connected sum of $M$ and $N$ along $L$.
Since $N$ is a closed manifold, $N/L$ is a compact pseudomanifold (with a CW structure) of dimension $n$ in the sense of [Dr1, Section 2.4]. The orientation sheaf ${\mathcal{O}}_{N/L}$ on $N/L$ gives a fundamental class $[N/L]$ that defines the Poincaré duality homomorphism $\text{PD}_{n}$ from Section 2.C in dimension $n$. When $N/L$ is orientable, $\text{PD}_{n}$ is an isomorphism for each coefficient system, see [Dr1, Theorem 2.4.1].
6.1 Theorem.
Let $M$ and $N$ be closed smooth orientable $n$-manifolds glued along $L$ (in the above sense) to form $M\#_{L}\hskip 0.56905ptN$ and let $N/L$ be a closed orientable $n$-psuedomanifold. If $N/L$ is aspherical, then $\operatorname{\mathsf{dcat}}(M\#_{L}\hskip 0.56905ptN)=n$. In particular, $M\#_{L}\hskip 0.56905ptN$ is essential.
Proof.
First, we note that since $N$ and $N/L$ are closed and orientable of dimension $n$, $N/L$ has non-trivial cohomology in degree $n$. Since $N/L$ is aspherical as well, we have $\operatorname{{\rm cd}}(\pi_{1}(N/L))\geq n$. Hence, $\operatorname{\mathsf{dcat}}(N/L)\geq n$ by Theorem 2.3. Due to Proposition 2.4, the fibration ${\mathcal{B}}_{n}(p):P(N/L)_{n}\to N/L$ does not admit a section. Hence, the primary obstruction $\kappa_{n}\in H^{n}(N/L;{\mathcal{F}})$ to a section of ${\mathcal{B}}_{n}(p)$ is non-zero. Let us define the collapsing map $\phi:M\#_{L}\hskip 0.56905ptN\to(M\#_{L}\hskip 0.56905ptN)/M=N/L$. By definition, $\phi^{*}(\kappa_{n})\in H^{n}(M\#_{L}\hskip 0.56905ptN,\phi^{*}({\mathcal{F}}))$ is the primary obstruction to a lift of $\phi$ with respect to ${\mathcal{B}}_{n}(p)$. Because $\dim(M\#_{L}\hskip 1.13809ptN)=n$, it is the only obstruction to such a lift of $\phi$. Due to orientability, $\phi$ sends ${\mathcal{O}}_{N/L}$ to the orientation sheaf ${\mathcal{O}}$ on $M\#_{L}\hskip 0.56905ptN$. Furthermore, $\text{deg}(\phi)=1$. Hence, in ${\mathcal{O}}_{N/L}\otimes{\mathcal{F}}$ coefficients, we get
@@ 6.A. Sums along submanifolds
due to Poincaré duality in $N/L$ in the top dimension. This means $\phi^{*}(\kappa_{n})\neq 0$ and so, a lift of $\phi$ does not exist with respect to ${\mathcal{B}}_{n}(p)$. Hence, using Propositions 2.5 (2) and 2.6, we get $n\leq\operatorname{\mathsf{dcat}}(\phi)\leq\operatorname{\mathsf{dcat}}(M\#_{L}\hskip 0.56905ptN)\leq\operatorname{\mathsf{cat}}(M\#_{L}\hskip 0.56905ptN)\leq\dim(M\#_{L}\hskip 0.56905ptN)=n$. Finally, we use Theorem 3.1 to conclude that $M\#_{L}\hskip 0.56905ptN$ is essential. ∎
@@
6.B. Doubles of punctured manifolds
We now consider a particular case of the generalized connected sum discussed above for compact manifolds. Let $M$ be a compact oriented $n$-manifold with boundary $\partial M\neq\emptyset$. We take a copy $\overline{M}$ of $M$ with reverse orientation and glue $M$ and $\overline{M}$ along $\partial M$ to form $\mathcal{D}M=M\#_{\partial M}\hskip 0.56905pt\overline{M}$. This is a closed oriented $n$-manifold which we call the double of $M$.
For any $k\geq 1$, let us call the process of the removal of $k$ number of small, disjoint open $n$-discs (or the interiors of closed $n$-discs) from a closed orientable $n$-manifold $M$ as a $k$-puncturing of $M$.
6.2 Proposition.
For $1\leq i\leq n$, let $k_{i}$, $r_{i}\geq 1$ and let $M_{i}$ be a closed orientable aspherical $r_{i}$-manifold. Let $X_{i}$ be obtained from $M_{i}$ after a $k_{i}$-puncturing of $M_{i}$. If $M=\prod_{i=1}^{n}X_{i}$, then $\operatorname{\mathsf{dcat}}(\mathcal{D}M)=\sum_{i=1}^{n}r_{i}\hskip 1.42262pt$ for the double $\mathcal{D}M$ of $M$. In particular, $\mathcal{D}M$ is essential.
Proof.
Let $Z=\prod_{i=1}^{n}M_{i}$ and $m=\sum_{i=1}^{n}r_{i}$. For each $i\leq n$, we can obtain $M_{i}$ back from $X_{i}$ by filling each of the $k_{i}$ punctures in $X_{i}$ by a suitable $r_{i}$-disc. Certainly, the inclusion $\mathcal{I}_{i}:X_{i}\hookrightarrow M_{i}$ is an embedding. So, there is a canonical map $\tau^{\prime}:M\to Z$. Let $q_{1}:M\to M/\partial M$ and $q_{2}:Z\to Z/\partial Z$ be the quotient maps that collapse the respective boundaries. Since $Z$ is closed, $q_{2}$ is the identity map on $Z$. Thus, $\tau^{\prime}$ produces a natural map $\tau:M/\partial M\to Z$. Let $\phi:\mathcal{D}M\to\mathcal{D}M/M=M/\partial M$ be the collapsing map such that $\text{deg}(\phi)=1$. Then, it follows that $(\tau\circ\phi)_{*}([\mathcal{D}M])=[Z]$. Since $Z$ is aspherical of dimension $m$, the primary obstruction $\kappa_{m}\in H^{m}(Z;{\mathcal{F}})$ to a section of the fibration ${\mathcal{B}}_{m}(p_{Z}):P(Z)_{m}\to Z$ is non-zero (see Section 3.A). Clearly, $(\tau\circ\phi)^{*}(\kappa_{m})\in H^{m}(\mathcal{D}M;(\tau\circ\phi)^{*}({\mathcal{F}}))$ is the primary obstruction to a lift of the map $\tau\circ\phi$ with respect to ${\mathcal{B}}_{m}(p_{Z})$, and since $\dim(\mathcal{D}M)=m$, it is the only obstruction to such a lift of $\tau\circ\phi$. We note that $\tau\circ\phi$ induces an isomorphism
@@ 6.B. Doubles of punctured manifolds
due to Poincaré duality in $Z$. Thus, $(\tau\circ\phi)^{*}(\kappa_{m})\neq 0$. Hence, we conclude that a lift of $\tau\circ\phi$ does not exist with respect to ${\mathcal{B}}_{m}(p_{Z})$. By Propositions 2.5 (2) and 2.6, we get $m\leq\operatorname{\mathsf{dcat}}(\tau\circ\phi)\leq\operatorname{\mathsf{dcat}}(\mathcal{D}M)\leq m$. So, $\operatorname{\mathsf{dcat}}(\mathcal{D}M)=\dim(\mathcal{D}M)=m$. Finally, we use Theorem 3.1 to conclude that $\mathcal{D}M$ is essential. ∎
@@
If $X_{i}$ is a punctured closed nilmanifold of dimension $r_{i}$ for $1\leq i\leq n$ and $M=\prod_{i=1}^{n}X_{i}$, then $\operatorname{\mathsf{dcat}}(\mathcal{D}M)=\sum_{i=1}^{n}r_{i}\hskip 1.42262pt$ by Proposition 6.2.
@@
In the notations of Proposition 6.2, more generally, we can take each $M_{i}$ to be a closed orientable essential $r_{i}$-manifold such that $B\pi_{1}(M_{i})$ is a closed orientable $r_{i}$-manifold for all $1\leq i\leq n$. Then we can proceed along the lines of Proposition 6.2 using a classifying map $g:Z\to\prod_{i=1}^{n}B\pi_{1}(M_{i})$ and some ideas from Theorem 4.1 to obtain $\operatorname{\mathsf{dcat}}(\mathcal{D}M)=\sum_{i=1}^{n}r_{i}$. So, in particular, we can take the non-aspherical symplectically aspherical branched covers $X^{n}$ from class $\mathscr{C}$ in Example 3.11, puncture them by deleting finitely many small, disjoint open $2n$-discs, take their product after puncturing, and conclude that the distributional LS-category of the double of that punctured product is equal to its dimension.
6.C. Geometrically decomposable $4$-manifolds
We follow Hillman [Hi] and say that a $4$-manifold $M$ is geometrically decomposable if it has a finite family $\mathcal{S}$ of disjoint $2$-sided hypersurfaces $S$ such that each component (called piece) of $M\setminus\cup S$ is geometric of finite volume.
@@ 6.C. Geometrically decomposable $4$-manifolds
Up to finite covers, the $4$-manifolds having $\mathbb{H}^{2}\times\mathbb{H}^{2}$ geometry are the products of two punctured surfaces with negative Euler characteristics. For $i=1,2$, let us fix some integers $g_{i}\geq 1$ and $k_{i}\geq 1$. Let $X_{i}$ be obtained from the surface $\Sigma_{g_{i}}$ after a $k_{i}$-puncturing of $\Sigma_{g_{i}}$. If $M=X_{1}\times X_{2}$, then its double $\mathcal{D}M=M\#_{\partial M}\hskip 0.56905pt\overline{M}$ is a closed orientable $4$-manifold belonging to class (3) described above and it is not aspherical. The simplest such example is the double of $T_{o}\times T_{o}$, where $T_{o}=T^{2}\setminus\text{int}(D^{2})$ is the once-punctured $2$-torus and $D^{2}$ denotes the closed $2$-disc.
@@
The boundary $\partial M$ of $M$ is a non-trivial graph $3$-manifold. When $k_{1}=k_{2}=1$, i.e., $X_{1}$ and $X_{2}$ are once-punctured surfaces, it is easy to describe $\partial M$ explicitly. Each $X_{i}$ has an infinite pipe, so let us visualize capping $X_{i}$ by a closed $2$-disc $D_{i}$ whose boundary $\partial D_{i}$ is the boundary of $X_{i}$ represented by a loop $t_{i}$. The boundary of $X_{i}\times\partial D_{i}$ is a $2$-torus $T_{i}$. Let $y_{i},z_{i}\in H_{1}(T_{i};{\mathbb{Z}})$ be the canonical generators for $i=1,2$. We choose a diffeomorphism $\phi:T_{1}\to T_{2}$ such that for the induced map $\phi_{*}$ in the first homology, $\phi_{*}(y_{1})=\pm z_{2}$ and $\phi_{*}(z_{1})=\pm y_{2}$. Then, $\partial M=((X_{1}\times\partial D_{1})\sqcup(X_{2}\times\partial D_{2}))/\sim$, where $x\sim\phi(x)$ for all $x\in T_{1}$. Basically, $\partial M$ is formed by identifying the loops $t_{1}$ and $t_{2}$ with the commutators of $X_{2}$ and $X_{1}$, respectively. If $a_{1},b_{1},\ldots,a_{g_{1}},b_{g_{1}}$ are loops in $X_{1}$ and $c_{1},d_{1},\ldots,c_{g_{2}},d_{g_{2}}$ are loops in $X_{2}$, then $\pi_{1}(\partial M)$ can be described as follows.
@@
As an application of Proposition 6.2, we obtain the following new computation for $\operatorname{\mathsf{dcat}}$ (and hence $\operatorname{\mathsf{cat}}$) of some closed non-aspherical geometrically decomposable $4$-manifolds belonging to class (3) of Theorem 6.5.
@@
If $M$ is the product of two punctured closed orientable surfaces of genera $\geq 1$, then $\operatorname{\mathsf{dcat}}(\mathcal{D}M)=4$ for the double $\mathcal{D}M$ of $M$.
@@
For $i=1,2$ and given $g_{i},k_{i}\geq 1$, we take $M_{i}=\Sigma_{g_{i}}$, $M=X_{1}\times X_{2}$, and $Z=\Sigma_{g_{1}}\times\Sigma_{g_{2}}$ in the proof of Proposition 6.2 and get $\operatorname{\mathsf{dcat}}(\mathcal{D}M)=4$. ∎

### v3

@@ 5.2. $4$-manifolds with free fundamental group
We will see more computations for $\operatorname{\mathsf{dcat}}$ of closed $4$-manifolds in Section 6.3.
5.3. An analog of Rudyak’s conjecture
@@
In the above setting, if $N={\mathbb{R}}P^{n}$ for odd $n$, then Question 5.10 is answered since $M$ is not contractible and thus, $\operatorname{\mathsf{dcat}}(M)\geq 1=\operatorname{\mathsf{dcat}}({\mathbb{R}}P^{n})$.
For $n\leq 2$, $\operatorname{\mathsf{dcat}}$ and $\operatorname{\mathsf{cat}}$ of closed orientable $n$-manifolds coincide, [DJ, Section 6.1]. So, Question 5.10 is answered affirmatively for $n\leq 2$ in view of Theorem 5.9.
@@
The proof of Theorem 5.9 for $n\in\{3,4\}$ uses the notions of category weight and detecting elements (see [R1, Section 2]). Unfortunately, we do not have useful analogous notions in the distributional theory yet, so we attempt to answer Question 5.10 affirmatively in some cases using our results from this paper.
@@
In case (1), we get $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)$ due to Theorem C. Similarly, in cases (2) and (3), $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)$ by our discussions in Section 5.2. In case (4), we apply Theorem B in dimensions $3$ and $4$ to get $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)$ again. Thus, in all these cases, Theorem 5.9 gives $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)\geq\operatorname{\mathsf{cat}}(N)\geq\operatorname{\mathsf{dcat}}(N)$. ∎
@@ 6. For generalized connected sums
6.1. Sums along submanifolds
Let $M$ and $N$ be two closed smooth orientable $n$-manifolds for $n\geq 3$. Let $S$ be a closed orientable manifold of dimension at most $n$. Suppose there are embeddings $i:S\to M$ and $j:S\to N$ with respective normal bundles $\mu$ and $\nu$ and an orientation-reversing bundle isomorphism $g:\mu\to\nu$. Such a map $g$ exists, in particular, when $\dim(S)=n-2$ and $e(\mu)=-e(\nu)$ for the Euler characteristics. Using smoothness, we can choose tubular neighborhoods $U\subset M$ and $V\subset N$ of $i(S)$ and $j(S)$, respectively, such that $U$ and $V$ are identified with the total spaces of $\mu$ and $\nu$, respectively. Then $g$ yields a diffeomorphism $h:U\to V$ that maps $i(S)$ to $j(S)$. This produces an orientation-reversing diffeomorphism $\phi:U-i(S)\to V-j(S)$, see [IKRT, Section 3]. Let us denote this space (up to homeomorphism) by $L$. We now take the disjoint union
$(M-i(S))\sqcup(N-j(S))$
and identify $U-i(S)$ with $V-j(S)$ using $\phi$ to obtain a closed orientable $n$-manifold $M\#_{L}\hskip 0.56905ptN$, which we call the generalized connected sum of $M$ and $N$ along $L$.
Since $N$ is a closed manifold, $N/L$ is a compact pseudomanifold (with a CW structure) of dimension $n$ in the sense of [Dr1, Section 2.4]. The orientation sheaf ${\mathcal{O}}_{N/L}$ on $N/L$ gives a fundamental class $[N/L]$ that defines the Poincaré duality homomorphism $\text{PD}_{n}$ from Section 2.3 in dimension $n$. When $N/L$ is orientable, $\text{PD}_{n}$ is an isomorphism for each coefficient system, see [Dr1, Theorem 2.4.1].
Proof of Theorem D.
First, we note that since $N$ and $N/L$ are closed and orientable of dimension $n$, $N/L$ has non-trivial cohomology in degree $n$. Since $N/L$ is aspherical as well, we have $\operatorname{{\rm cd}}(\pi_{1}(N/L))\geq n$. Hence, $\operatorname{\mathsf{dcat}}(N/L)\geq n$ by Theorem 2.3. Due to Proposition 2.4, the fibration ${\mathcal{B}}_{n}(p):P(N/L)_{n}\to N/L$ does not admit a section. Hence, the primary obstruction $\kappa_{n}\in H^{n}(N/L;{\mathcal{F}})$ to a section of ${\mathcal{B}}_{n}(p)$ is non-zero. Let us define the collapsing map
$\phi:M\#_{L}\hskip 0.56905ptN\to(M\#_{L}\hskip 0.56905ptN)/M=N/L.$
By definition, $\phi^{*}(\kappa_{n})\in H^{n}(M\#_{L}\hskip 0.56905ptN,\phi^{*}({\mathcal{F}}))$ is the primary obstruction to a lift of $\phi$ with respect to ${\mathcal{B}}_{n}(p)$. Because $\dim(M\#_{L}\hskip 1.13809ptN)=n$, it is the only obstruction to such a lift of $\phi$. Due to orientability, $\phi$ sends ${\mathcal{O}}_{N/L}$ to the orientation sheaf ${\mathcal{O}}$ on $M\#_{L}\hskip 0.56905ptN$. Furthermore, $\text{deg}(\phi)=1$. Hence, in ${\mathcal{O}}_{N/L}\otimes{\mathcal{F}}$ coefficients, we get
@@ 6.1. Sums along submanifolds
due to Poincaré duality in $N/L$ in the top dimension. This means $\phi^{*}(\kappa_{n})\neq 0$ and so, a lift of $\phi$ does not exist with respect to ${\mathcal{B}}_{n}(p)$. Hence, using Propositions 2.5 (2) and 2.6, we get the inequality
$n\leq\operatorname{\mathsf{dcat}}(\phi)\leq\operatorname{\mathsf{dcat}}(M\#_{L}\hskip 0.56905ptN)\leq\operatorname{\mathsf{cat}}(M\#_{L}\hskip 0.56905ptN)\leq\dim(M\#_{L}\hskip 0.56905ptN)=n.$
@@
6.1 Remark.
It is not known if the connected sum formula [DS] for classical LS-category ($\operatorname{\mathsf{cat}}$) works for certain generalized connected sums as well, such as those studied in this section. Hence, Theorem D potentially provides new computations of $\operatorname{\mathsf{cat}}$ for some special fiber sums.
6.2. Doubles of punctured manifolds
We now consider a particular case of the generalized connected sum discussed above for compact manifolds. Let $M$ be a compact oriented $n$-manifold with boundary $\partial M\neq\emptyset$. We take a copy $\overline{M}$ of $M$ with reverse orientation and glue $M$ and $\overline{M}$ along $\partial M$ to form
$\mathcal{D}M=M\#_{\partial M}\hskip 0.56905pt\overline{M}.$
This is a closed oriented $n$-manifold which we call the double of $M$.
Recall that any $k\geq 1$, a $k$-puncturing of a closed orientable $n$-manifold $M$ refers to the process of the removal of $k$ number of small, disjoint open $n$-discs (or the interiors of closed $n$-discs) from $M$.
Proof of Theorem E.
Let $Z=\prod_{i=1}^{n}M_{i}$ and $m=\sum_{i=1}^{n}r_{i}$. For each $i\leq n$, we can obtain $M_{i}$ back from $X_{i}$ by filling each of the $k_{i}$ punctures in $X_{i}$ by a suitable $r_{i}$-disc. Certainly, the inclusion $\mathcal{I}_{i}:X_{i}\hookrightarrow M_{i}$ is an embedding. So, there is a canonical map $\tau^{\prime}:M\to Z$. Let $q_{1}:M\to M/\partial M$ and $q_{2}:Z\to Z/\partial Z$ be the quotient maps that collapse the respective boundaries. Since $Z$ is closed, $q_{2}$ is the identity map on $Z$. Thus, $\tau^{\prime}$ produces a natural map $\tau:M/\partial M\to Z$. Let
$\phi:\mathcal{D}M\to\mathcal{D}M/M=M/\partial M$
be the collapsing map with $\text{deg}(\phi)=1$. Then, it follows that $(\tau\circ\phi)_{*}([\mathcal{D}M])=[Z]$. Since $Z$ is aspherical of dimension $m$, the primary obstruction $\kappa_{m}\in H^{m}(Z;{\mathcal{F}})$ to a section of the fibration ${\mathcal{B}}_{m}(p_{Z}):P(Z)_{m}\to Z$ is non-zero (see Section 3.1). Clearly,
$(\tau\circ\phi)^{*}(\kappa_{m})\in H^{m}(\mathcal{D}M;(\tau\circ\phi)^{*}({\mathcal{F}}))$
is the primary obstruction to a lift of the map $\tau\circ\phi$ with respect to ${\mathcal{B}}_{m}(p_{Z})$, and since $\dim(\mathcal{D}M)=m$, it is the only obstruction to such a lift of $\tau\circ\phi$. We note that $\tau\circ\phi$ induces an isomorphism
@@ 6.2. Doubles of punctured manifolds
due to Poincaré duality in $Z$. Thus, $(\tau\circ\phi)^{*}(\kappa_{m})\neq 0$. Hence, we conclude that a lift of $\tau\circ\phi$ does not exist with respect to ${\mathcal{B}}_{m}(p_{Z})$. Due to Propositions 2.5 (2) and 2.6, we have that
$m\leq\operatorname{\mathsf{dcat}}(\tau\circ\phi)\leq\operatorname{\mathsf{dcat}}(\mathcal{D}M)\leq m.$
Therefore, $\operatorname{\mathsf{dcat}}(\mathcal{D}M)=\dim(\mathcal{D}M)=m$. ∎
@@
If $X_{i}$ is a punctured closed nilmanifold of dimension $r_{i}$ for $1\leq i\leq n$ and $M=\prod_{i=1}^{n}X_{i}$, then $\operatorname{\mathsf{dcat}}(\mathcal{D}M)=\sum_{i=1}^{n}r_{i}\hskip 1.42262pt$ by Theorem E.
@@
In the notations of Theorem E, more generally, we can take each $M_{i}$ to be a closed orientable essential $r_{i}$-manifold such that $B\pi_{1}(M_{i})$ is a closed orientable $r_{i}$-manifold for all $1\leq i\leq n$. Then we can proceed along the lines of Theorem E using a classifying map $g:Z\to\prod_{i=1}^{n}B\pi_{1}(M_{i})$ and some ideas from Theorem B to obtain $\operatorname{\mathsf{dcat}}(\mathcal{D}M)=\sum_{i=1}^{n}r_{i}$. So, in particular, we can take the non-aspherical symplectically aspherical branched covers $X^{n}$ from class $\mathscr{C}$ in Example 3.10, puncture them by deleting finitely many small, disjoint open $2n$-discs, take their product after puncturing, and conclude that the distributional category of the double of that punctured product is equal to its dimension.
We mention that we have not been able to find results in the existing literature showing that the above double manifolds are essential. So, we believe, Theorem E gives new computation of the classical LS-category of these manifolds using purely obstruction-theoretic techniques.
6.3. Geometrically decomposable $4$-manifolds
We follow Hillman [Hi] and say that a $4$-manifold $M$ is geometrically decomposable if it has a finite family $\mathcal{S}$ of disjoint $2$-sided hypersurfaces $S$ such that each component (called a piece) of $M\setminus\cup S$ is geometric of finite volume.
@@ 6.3. Geometrically decomposable $4$-manifolds
Up to finite covers, the $4$-manifolds having $\mathbb{H}^{2}\times\mathbb{H}^{2}$ geometry are the products of two punctured surfaces with negative Euler characteristics. For $i=1,2$, let us fix some integers $g_{i}\geq 1$ and $k_{i}\geq 1$. Let $X_{i}$ be obtained from the surface $\Sigma_{g_{i}}$ after a $k_{i}$-puncturing of $\Sigma_{g_{i}}$. If $M=X_{1}\times X_{2}$, then its double
$\mathcal{D}M=M\#_{\partial M}\hskip 0.56905pt\overline{M}$
is a closed orientable $4$-manifold belonging to class (3) described above and it is not aspherical. The simplest such example is the double of $T_{o}\times T_{o}$, where $T_{o}=T^{2}\setminus\text{int}(D^{2})$ is the once-punctured $2$-torus and $D^{2}$ denotes the closed $2$-disc.
@@
The boundary $\partial M$ of $M$ is a non-trivial graph $3$-manifold. When $k_{1}=k_{2}=1$, i.e., $X_{1}$ and $X_{2}$ are once-punctured surfaces, it is easy to describe $\partial M$ explicitly. Each $X_{i}$ has an infinite pipe, so let us visualize capping $X_{i}$ by a closed $2$-disc $D_{i}$ whose boundary $\partial D_{i}$ is the boundary of $X_{i}$ represented by a loop $t_{i}$. The boundary of $X_{i}\times\partial D_{i}$ is a $2$-torus $T_{i}$. Let $y_{i},z_{i}\in H_{1}(T_{i};{\mathbb{Z}})$ be the canonical generators for $i=1,2$. We choose a diffeomorphism $\phi:T_{1}\to T_{2}$ such that for the induced map $\phi_{*}$ in the first homology, $\phi_{*}(y_{1})=\pm z_{2}$ and $\phi_{*}(z_{1})=\pm y_{2}$. Then,
$\partial M=((X_{1}\times\partial D_{1})\sqcup(X_{2}\times\partial D_{2}))/\sim,$
where $x\sim\phi(x)$ for all $x\in T_{1}$. Basically, $\partial M$ is formed by identifying the loops $t_{1}$ and $t_{2}$ with the commutators of $X_{2}$ and $X_{1}$, respectively. If $a_{1},b_{1},\ldots,a_{g_{1}},b_{g_{1}}$ are loops in $X_{1}$ and $c_{1},d_{1},\ldots,c_{g_{2}},d_{g_{2}}$ are loops in $X_{2}$, then $\pi_{1}(\partial M)$ can be described as follows.
@@
As an application of Theorem E, we obtain the following new computation for $\operatorname{\mathsf{dcat}}$ (and hence $\operatorname{\mathsf{cat}}$) of some closed non-aspherical geometrically decomposable $4$-manifolds belonging to class (3) of Theorem 6.4.
@@
If $M$ is the product of two finitely punctured closed orientable surfaces of genera $\geq 1$, then $\operatorname{\mathsf{dcat}}(\mathcal{D}M)=4$ for the double $\mathcal{D}M$ of $M$.
@@
For $i=1,2$ and given $g_{i},k_{i}\geq 1$, we take $M_{i}=\Sigma_{g_{i}}$, $M=X_{1}\times X_{2}$, and $Z=\Sigma_{g_{1}}\times\Sigma_{g_{2}}$ in the proof of Theorem E and get $\operatorname{\mathsf{dcat}}(\mathcal{D}M)=4$. ∎
It seems that a proof of the essentiality of the closed non-aspherical geometrically decomposable $4$-manifolds belonging to class (3) of Theorem 6.4 is missing from the literature. Our Corollary 6.6 fills this apparent gap.